### v1

# Aligning Translation-Specific Understanding to General Understanding in Large Language Models

Abstract
Although large language models (LLMs) have shown surprising language understanding and generation capabilities, they have yet to gain a revolutionary advancement in the field of machine translation. One potential cause of the limited performance is the misalignment between the translation-specific understanding and general understanding inside LLMs. To align the translation-specific understanding to the general one, we propose a novel translation process xIoD (Cross-Lingual Interpretation of Difficult words), explicitly incorporating the general understanding on the content incurring inconsistent understanding to guide the translation. Specifically, xIoD performs the cross-lingual interpretation for the difficult-to-translate words and enhances the translation with the generated interpretations. Furthermore, we reframe the external tools of QE to tackle the challenges of xIoD in the detection of difficult words and the generation of helpful interpretations. We conduct experiments on the self-constructed benchmark Challenge-MT, which includes cases in which multiple SOTA translation systems consistently underperform. Experimental results show the effectiveness of our xIoD, which improves up to +3.85 COMET. ††footnotetext: 🖂 means corresponding author.
Aligning Translation-Specific Understanding to General Understanding in Large Language Models
Yichong Huang, Xiaocheng Feng, Baohang Li, Chengpeng Fu, Wenshuai Huo, Ting Liu, Bing Qin Harbin Institute of Technology Peng Cheng Laboratory {ychuang,xcfeng,baohangli,cpfu,wshuo,tliu,qinb}@ir.hit.edu.cn
1 Introduction
Recently, with the scaling of model capacities and pre-training data volume, large language models (LLMs) have demonstrated remarkable language understanding and generation, paving the way for a higher level of performance in machine translation Zhao et al. (2023); OpenAI (2023); Jiang et al. (2023); Workshop (2023). However, existing research reports that LLMs have yet to achieve as significant advances in machine translation as they have achieved in other natural language processing fields Hendy et al. (2023); Zhang et al. (2023a); Jiao et al. (2023b); Zhu et al. (2023b).
One potential reason leading to the limited translation performance of LLMs is the misalignment between the general understanding and translation-specific understanding inside LLMs. This claim is based on our discovery that LLMs understand many concepts accurately when explaining directly but often misunderstand these concepts in translation requests, which is illustrated in Fig 1. This paper refers to these concepts as language models’ generalization failures on translation.
In this work, we propose a novel translation process xIoD, which explicitly incorporates the general understanding to guide the translation during inference. Specifically, xIoD first detects the difficult-to-translation words in the source sentence, which could cover the generalization failures intuitively. Next, the LLM is prompted to interpret each difficult word with the target language, i.e., cross-lingual interpretation of difficult words, unleashing the powerful general understanding and aligning this understanding into the target language space. After that, xIoD conducts translation under the guidance of these interpretations. Unlike the CoT-based process mimicking the junior translator to perform word-for-word translation Peng et al. (2023), xIoD is more in accordance with the senior translator owning a higher performance frontier. However, xIoD faces two challenges: (1) LLMs may struggle with the accurate detection of difficult words due to their limitation in self-knowledge Yin et al. (2023), and (2) LLMs are prone to errors and hallucinations in the interpretations Huang et al. (2023a, b), biasing the translation from the original semantics. Therefore, we reframe the external tool of token-level QE Rei et al. (2023) to enhance the detection of difficult words and design a strategy of interpretations quality control to filter hallucinated interpretations based on sentence-level QE Rei et al. (2020).
Towards better identifying the limitations of current MT systems, we proposed the benchmark Challenge-MT via collecting the difficult cases that multiple SOTA systems consistently underperform on multi-year WMT datasets. All experiments are conducted under the few-shot setting, where the demonstrations are constructed automatically in a post-explanation manner. Experimental results on six translation directions (Chinese, Estonian, and Icelandic to/from English) demonstrate the effectiveness of xIoD, which improves up to +3.85 COMET.
2 Background
2.1 LLM-based MT
Considering the translation from source language to target language , LLM-based machine translation converts the source sentence to an instruction using a translation-specific template and generates the translation by feeding the instruction to the LLM . To make the LLM better follow the instruction, the in-context learning (ICL) strategy Brown et al. (2020); Dong et al. (2023) injects a few examples/demonstrations of translation into the instruction, which is shown as:
Formally, the LLM-based MT generates the translation with ICL as:
where is the demonstrations set of translation.
2.2 QE for MT
Quality estimation (QE) for machine translation, i.e., reference-free MT evaluation, aims to predict the quality of the given translation only according to the source sentence, which has shown auspicious correlations with human judgments Rei et al. (2020, 2021). Given a source sentence and the translation , the QE score is denoted as .
Thanks to the recent advance in the interpretability of neural MT metrics Rei et al. (2023), token-level QE is proposed to score the error degree of the given translation span by calculating the misalignment of this span against the source sentence. Given a source sentence and the candidate translation , token-level QE annotates the error degree of the specific span in the translation, i.e., where .
3 Approach: xIoD
In this section, we first introduce our translation framework xIoD (§3.1). Specifically, xIoD consists of three components: difficult word detection (§3.2), cross-lingual interpretation (§3.3), and interpretation quality control (§3.4). To make the LLM follow the procedure of each component as expected, we adopt the in-context learning strategy and design an automatic method for constructing demonstrations of xIoD (§3.5).
3.1 Framework
The progress of xIoD is illustrated in Figure 2. Given the source sentence, xIoD first detects the difficult words or phrases in the source sentence. Once the difficult words are identified, xIoD requests the LLM to interpret each difficult word with the target language, unleashing the powerful understanding capability inside the LLM and aligning these understanding into the target language. Finally, to avoid the negative effect of incorrect and useless interpretations, xIoD removes the negative interpretations through the interpretation quality control and outputs the final translation guided by the helpful interpretations. We describe each step in the later.
3.2 Step-1: Difficult Word Detection
Correct understanding of the difficult words determines largely the quality of the final translation. To make the LLM comprehend these difficult words, it is necessary to recognize them at first. However, finding the translation-specific difficult words remains challenging Lim et al. (2023); Sun (2015). To tackle this challenge, we first conduct a preliminary translation for the given source sentence and then extract the mistranslated words in the source sentence as the difficult words. Concretely, we invent xIoD-I to do this using solely the Intrinsic knowledge of LLMs at first.
xIoD-I.
Given source sentence , xIoD-I first obtains the preliminary translation (also known as draft translation) by prompting the LLM to translate with the in-context learning strategy, which is shown in Eq. (1). Next, the LLM is requested to output the difficult words based on the source sentence and the preliminary translation:
xIoD-I obtains the difficult word list via performing greedy decoding on the LLM:
where is the LLM, which is prompted with demonstrations of difficult word detection .
xIoD-E.
It is frequently observed that the LLM fails to recognize the mistranslated words. Therefore, we devise xIoD-E to boost the detection with the external tool. First, xIoD-E requests the LLM with the same prompt as xIoD-I while performing the temperature sampling for times (). Next, the union of all sampling results is taken as the candidate set of difficult words :
where is the sampling temperature††https://platform.openai.com/docs/api-reference/chat/create#chat-create-temperature, which is set to to capture more candidates.
Finally, xIoD-E annotates each candidate word with its degree of misalignment with respect to the draft translation, which reflects the translation-specific difficulty. To implement this function, we adopt an external tool of token-level QE . As shown in §2.2, token-level QE is originally used to score the mistranslation degree of the given translation span with respect to the source sentence, i.e., where . Differently, we utilize this tool in a dual manner. That is, we use to annotate the misalignment degree of the given source span with respect to the translation, i.e., where . Formally, the misalignment score of each difficult word candidate is calculated as:
Then, xIoD-E selects candidates with misalignment score , where is the hyperparameter named the difficulty threshold. We refer to this procedure as the difficulty-aware selection in Fig 2.
3.3 Step-2: Cross-Lingual Interpretation
After the difficult words in the source sentence are detected, xIoD lets the LLM generate the interpretation of each difficult word via requesting with the prompt:
Through access to the LLM, the interpretation set is obtained:
where , which is the demonstrations of the cross-lingual interpretation.
Prob and cons.
Through the generated interpretations, xIoD enhances the translation with LLMs‘ general understanding capability. However, LLMs may generate incorrect or hallucinated interpretations sometimes (e.g., the interpretation of ”崩塌” in Fig 2), which biases the resulting translation from the original semantics. Besides, helpless interpretations that can not provide useful information also pose a risk of disturbing the translation process.
3.4 Step-3: Interpretation Quality Control
To overcome the potential negative effect of the generated interpretations, xIoD removes the incorrect and useless interpretations through the interpretation quality control (IQC) and outputs the final translation guided by the helpful interpretations.
Concretely, given a set of interpretations , xIoD ablates each interpretation sequentially and uses the remaining interpretations to guide the translation. The interpretation-guided translation is implemented in a fashion of refinement:
Formally, the translation is obtained as:
where , which is the demonstrations of interpretation-guide translation.
Once the better translation performance is achieved by ablation, which is measured by the QE††We use wmt21-comet-qe-da as the QE scorer. tool due to the unavailable access to the reference translation, the interpretation is removed from and the current translation is taken as the best translation. We detail this process in Alg. 1.
3.5 Demonstrations Synthesis for xIoD
To make the LLM follow the procedure of xIoD as expected, we adopt the ICL strategy. Common practice constructs the demonstrations manually for ICL, necessitating human translators proficient in language pairs for languages. To overcome this considerable cost, we devise a method for synthesizing high-quality demonstrations of xIoD based on the parallel data.
Inspired by the idea of Auto-CoT Zhang et al. (2023b), we utilize LLM to generate the difficult words and corresponding interpretations based on the given bilingual sentence pair :
Then, the response is parsed via regular expression to extract the difficult words and interpretations . Next, we remove the noisy interpretations through a process similar to IQC (Alg. 1). The only difference is that the QE metric is replaced with the reference-based COMET Rei et al. (2020) due to the available access to the reference translation. Finally, the generated difficult words and interpretations can be assembled with the source and target sentence as demonstrations for each step of xIoD.
| Methods | EnZh | ZhEn | EnEt | EtEn | EnIs | IsEn | Average | |||||||
| COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | |
| Existing Systems | ||||||||||||||
| 74.85 | 1.87 | 68.21 | -5.97 | 79.11 | 6.02 | 78.83 | 5.57 | 76.17 | 0.56 | 78.70 | 3.14 | 75.98 | 1.87 | |
| NLLB | 68.77 | -2.74 | 60.09 | -11.27 | 74.20 | 2.40 | 74.35 | 3.57 | 69.37 | -5.28 | 72.55 | 0.90 | 69.89 | -2.07 |
| GPT-4 | 76.15 | 3.00 | 70.77 | -1.40 | 80.25 | 6.28 | 77.83 | 5.35 | 77.33 | 1.61 | 79.39 | 3.54 | 76.95 | 3.06 |
| Baselines | ||||||||||||||
| Zero-shot | 74.89 | 1.76 | 71.27 | -1.34 | 80.67 | 7.55 | 74.93 | 5.45 | 71.17 | -2.73 | 76.22 | 2.70 | 74.86 | 2.23 |
| ICL | 75.47 | 2.31 | 72.22 | -0.73 | 80.9 | 9.37 | 79.40 | 6.56 | 73.19 | -1.17 | 77.52 | 3.55 | 76.45 | 3.32 |
| +CoT | 73.85 | 0.86 | 71.35 | -1.00 | 78.03 | 5.24 | 76.78 | 4.77 | 69.72 | -2.96 | 76.55 | 2.96 | 74.38 | 1.65 |
| +Topic | 75.83 | 2.40 | 72.46 | -0.25 | 80.98 | 7.19 | 79.20 | 6.59 | 72.77 | -1.58 | 76.49 | 2.80 | 76.29 | 2.86 |
| +Keywords | 73.93 | 1.02 | 71.22 | -1.62 | 78.63 | 5.87 | 77.79 | 5.48 | 70.33 | -3.40 | 74.55 | 1.48 | 74.41 | 1.47 |
| +SimDems | 75.22 | 2.06 | 72.20 | -0.60 | 81.24 | 7.90 | 79.11 | 6.76 | 72.70 | -1.52 | 76.78 | 3.01 | 76.21 | 2.94 |
| Ours | ||||||||||||||
| xIoD-I | 76.92 | 4.07 | 72.94 | 0.45 | 82.92 | 9.68 | 79.96 | 7.19 | 76.64 | 2.98 | 78.45 | 4.65 | 77.97 | 4.84 |
| xIoD-E | 77.57 | 4.61 | 73.23 | 0.32 | 83.07 | 10.39 | 80.01 | 7.81 | 77.04 | 3.03 | 78.70 | 4.82 | 78.27 | 5.16 |
4 Testbed: Challenge-MT dataset
In this work, we propose a benchmark Challenge-MT, which consists of difficult translation samples, for the following reasons: 1) MT has witnessed human-like performance on many languages, which appeals to a more challenging benchmark to detect the flaws of SOTA systems, and 2) a benchmark containing more instances that models underperform helps us to analyze the issue of the understanding misalignment in LLMs effectively.
Challenge-MT is constructed by collecting the most challenge subset of the widely used WMT datasets, involving six translation directions of Chinese (zh), Estonian (et), and Icelandic (is) to/from English (en). Specifically, we first evaluate three SOTA MT systems (Google Translate, ChatGPT (gpt-3.5-turbo), and NLLB NLLB Team et al. (2022)) on the multi-year WMT datasets. Due to the poor performance of NLLB on the zhen translation, we additionally train a zhen translation model based on the DeltaLM Ma et al. (2021) on the parallel corpus from OPUS††https://opus.nlpl.eu/. Next, The generated translations are scored with the COMET metric, and the of instances with the lowest score for each system are extracted as its difficult samples set. We vary the value of across different language pairs to ensure an appropriate scale for each difficult sample set. Finally, the intersection of all systems’ difficult sample sets is taken as the Challenge-MT benchmark. We equally split this dataset into the validation set and the test set.
As the results demonstrated in Fig 3, SOTA MT systems show extremely poor performance on the Challenge-MT benchmark, which is 10 COMET scores lower than the complete set in most translation directions. To understand the cause of the low performance on Challenge-MT, we conduct a multi-aspect comparison for the complete set and the Challenge-MT subset in Appendix A, which shows that the samples are longer and have higher perplexity in the Challenge-MT. These features indicate that sentences that are longer and harder to understand are more difficult to translate.
5 Experiments
5.1 Experimental Setup
Comparative Methods.
We verify the effectiveness of our xIoD on the LLM GPT-3.5-turbo for its promising capability in following complicated instructions. Demonstrations of xIoD are gained by performing our automatic method (§3.5) on the validation set of Challenge-MT. We compare xIoD with the following methods:
Zero-shot, which asks the LLM to translate the source sentence directly.
ICL (in-context learning), enhancing the translation with randomly selected exemplars from the validation set.
CoT Wei et al. (2022), encouraging the LLM to resolve the problem step by step. In this work, we re-implement CoT by prompting the LLM to translate the source sentence step by step.
MAPS He et al. (2023), incorporating the knowledge of keywords, topic words, and demonstrations similar to the given source sentence to enhance the translation process, respectively.
Commercial and open-source systems. We also report the performance of Google Translate, NLLB (in zhen translation, we replace NLLB with our trained MT model based on DeltaLM), and zero-shot translation based on GPT4 (GPT-4-turbo).
For xIoD and other ICL-based methods, we select =8 demonstrations (i.e., 8-shot) to achieve a strong baseline performance. More details of re-implementing the baselines under the few-shot setting are illustrated in Appendix B
Metrics.
Following previous research of LLM-based MT Garcia et al. (2023); Chen et al. (2023a), we adopt COMET Rei et al. (2020) as the evaluation metric as its higher correlation with human judgment than BLEU Papineni et al. (2002). Besides, we report the QE score to alleviate the reference bias Freitag et al. (2020) of the reference-based metrics.
5.2 Results
| Methods | EnIs | IsEn | ||
| COMET | QE | COMET | QE | |
| Zero-shot | 77.33 | 1.61 | 79.39 | 3.54 |
| ICL | 80.1 | 3.76 | 81.02 | 4.33 |
| xIoD-E | 81.7 | 7.00 | 81.21 | 5.75 |
The main results are illustrated in Table 1. From the results, we have drawn the following observations:
(1) xIoD achieves significant improvements.
Both xIoD-I and xIoD-E outperform the baseline ICL significantly. Specifically, xIoD-E achieves significant improvements of +1.72 and +3.31 COMET over the baseline ICL and the Zero-shot method in average, respectively. In the low-source EnIs translation, xIoD-E improves the ICL by +3.45 COMET and the Zero-shot by +5.47 COMET. These improvements demonstrate that xIoD elicits the translation ability largely through aligning the translation-specific understanding to the general one.
(2) xIoD achieves state-of-the-art performance.
xIoD achieves the highest scores in EnZh and EnEt translation in terms of COMET and achieves the highest scores across all language pairs in terms of QE. However, due to the stronger multilingual capabilities of GPT-4 over GPT-3.5, the results of xIoD based on GPT-3.5 are lower than the results of the zero-shot method based on GPT-4 in EnIs translation. To verify the effectiveness of xIoD on LLMs with stronger multilingual capabilities, we implement xIoD based on GPT-4 in EnIs translation, as shown in Table 2. The improvements achieved by xIoD suggest that LLMs with stronger multilingual capabilities also suffer from the issue of understanding misalignment.
(3) CoT works poorly in machine translation.
As illustrated in Table 1, CoT incurs a dramatic performance drop on the baseline ICL. To investigate this phenomenon, we have conducted case analysis and observed that the translations generated by the CoT are extremely wordy and unfluent. Our finding aligns with those reported in Peng et al. (2023). This consistency suggests that the sequentially step-by-step paradigm may not be effectively applicable to the translation task.
(4) xIoD surpasses previous translation strategies.
The results show that incorporating the analysis of keywords and topics has not achieved a consistent improvement as xIoD. We conjecture the reason is that it is the difficult-to-translate words that lead to the performance bottleneck of MT due to the long-tail distribution of knowledge Kandpal et al. (2023); Feng et al. (2023). Besides, we also follow He et al. (2023) to experiment under the rerank setting as shown in Appendix.C, which demonstrates the effectiveness of our method further.
(5) xIoD-E achieves a further improvement over xIoD-I.
Specifically, xIoD-E outperforms xIoD-I by +0.3 COMET on average. This modest improvement demonstrates that the LLM (GPT-3.5-turbo) has reached an acceptable level of detecting difficult words, and the external tool of token-level QE could enhance its capability further.
6 Analysis
6.1 Ablation Study
| Methods | EnZh | ZhEn | ||
| COMET | COMET | |||
| xIoD-E | 77.57 | – | 73.23 | – |
| w/o. Draft | 76.94 | -0.63 | 72.68 | -0.55 |
| w/o. IQC | 76.54 | -1.03 | 72.91 | -0.32 |
| xIoD-I | 76.92 | – | 72.94 | – |
| w/o. Draft | 76.68 | -0.24 | 72.78 | -0.16 |
| w/o. IQC | 76.45 | -0.47 | 72.59 | -0.35 |
xIoD introduces the processes of (1) draft translation to precisely detect the difficult words and (2) IQC to improves the correctness of interpretations. To clearly elucidate the contribution of these two components, we conduct an ablation study in Table 3. Specifically, we analyze the effect of the draft translation by asking the LLM to detect difficult words directly without the draft translation. The impact of IQC is analyzed by evaluating the performance of the generated translations guided by the original noisy interpretations (i.e., without the processing of IQC). The results show that removing either component leads to performance drops, and IQC plays a more important role in xIoD. Specifically, the improvement of xIoD is halved when ablating the IQC on the EnZh translation.
6.2 Analysis of Difficult Word Detection
To offer an in-depth insight into the process of difficult word detection, we illustrate the relation between the number of difficult words interpreted and the resulting performance by adjusting the value of the difficulty threshold (), which is shown in Fig 4. Concretely, a smaller value of allows more difficult words to be interpreted. From the results, we have the following observations:
Increasing the number of interpretations does not necessarily lead to performance improvements, but increasing high-quality ones can.
Specifically, without controlling the quality of the interpretations (i.e., w/o. IQC), increasing the number of interpretations (the green lines) yields unpredictable performance changes (as shown by the green bins), as introducing either valuable information or noise. Fortunately, with IQC filtering negative interpretations, increasing the number of interpretations (the blue lines) leads to constant improvements (as the blue bins show).
Interpreting words that are more difficult brings larger improvements.
Specifically, in the EnZh translation, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to (). However, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to ().
It should be noted that interpreting more words incurs more inference costs. Therefore, a modest value of (i.e., ) is recommended to reach a compromise between efficiency and performance of xIoD.
6.3 Analysis of Interpretation Generation
Languages of interpretations.
Given a difficult word, xIoD generates the corresponding interpretation with the target language (i.e., cross-lingual interpretation), which implicitly comprises two stages: (1) generating the interpretation in the source language and (2) translating the interpretation into the target language. Compared with conducting these two stages explicitly, xIoD is more efficient and avoids error accumulation, which is illustrated in Fig 5. As demonstrated, interpretations in the target language (the blue bins) are more beneficial than the ones in the source language (the purple bins) owing to aligning the general understanding into the target language space, which could provide more benefits for translation. And the implicit two-stage process (the blue bins) is better than the explicit one (the green bins).
7 Related Work
Evaluation of LLMs’ translation capabilities.
With the remarkable progress of LLMs, researchers have assessed their translation abilities in various aspects. Zhang et al. (2023a); Vilar et al. (2023); Garcia et al. (2023); Bawden and Yvon (2023) have first identified best practice of LLM-based MT in terms of the prompt template and examples selection. Next, the evaluation is extended across more domains Hendy et al. (2023), more languages Zhu et al. (2023a), and document-level translation Hendy et al. (2023); Wang et al. (2023). Other lines of work have performed in-depth assessments on the important attributes beyond accuracy, like literalness Raunak et al. (2023) and culture awareness Yao et al. (2023). As existing studies have shown that LLMs have achieved promising performance especially on high-resource languages, our work turns out to benchmark them on hard instances, which are longer and more difficult to translate, towards detecting more underlying issues.
Translation-oriented fine-tuning.
To reach a new SOTA performance of MT, recent works have explored fine-tuning the LLM with parallel data. Specifically, Jiao et al. (2023a); Chen et al. (2023b) prove the effectiveness of incorporating both high-quality translations and human-annotated low-quality translations during fine-tuning. Alves et al. (2023) incorporate few-shot examples in the stage of fine-tuning to obtain both advantages of ICL and fine-tuning. Xu et al. (2023) propose the two-stage paradigm: initial fine-tuning on monolingual data followed by subsequent fine-tuning on a small set of high-quality parallel data. Zhu et al. (2023b) empower the LLM by tuning it with translation task data and cross-lingual general task data. Differently, our work aims to improve the translation performance of LLM without parameter updating, which avoids the potential risk of catastrophic forgetting on tasks beyond translation.
LLM-based translation strategies.
Gao et al. (2023) introduce the domain and POS information in the inference stage. Lu et al. (2023) obtain the multilingual translations of keywords in the source sentence via the translator NLLB to augment the LLM, which improves the translation of low-resource languages while hurting the performance of high-source languages. Chen et al. (2023a) demonstrate that iterative refinement reduces translationese significantly. He et al. (2023) incorporate the knowledge of keywords, topics, and demonstrations similar to the given source sentence to enhance the translation process, respectively. Finally, a rerank-like knowledge selection is used to combine all candidate translations. However, there is no significant improvement to be observed when solely utilizing each single type of knowledge. While our xIoD could fall into the category of ”translation strategy,” we are the first to address the misalignment between the translation-specific understanding and the general understanding in LLM and achieve significant improvement in both low-resource languages and high-resource ones. In the view of methodology, different from previous works that utilize the intrinsic knowledge of LLM, xIoD focuses on dealing with the difficult-to-translate words instead of the keywords for the reason that we argue the difficult-to-translate words lead to the performance bottleneck due to the long-tail distribution of knowledge.
8 Conclusion
In this work, we take the first step in resolving the misalignment between the translation-specific understanding and the general understanding. To this effect, we propose a novel translation process xIoD, incorporating the interpretations of the difficult-to-translate words, which carry the general understanding of LLM, to guide the translation. Furthermore, we utilize the token-level QE to enhance the detection of difficult words and the sentence-level QE to remove harmful interpretations. Experimental results on the proposed Challenge-MT benchmark illustrate the effectiveness of our method.
9 Limitations
Even though xIoD elicits the translation abilities of LLMs via unleashing the general understanding (intrinsic knowledge) of LLMs, they still struggle to translate concepts that require the incorporation of extrinsic knowledge, such as the translation of neologisms. However, Our approach lays the foundation for researching when and how to incorporate external knowledge.
Appendix A Fine-grained Statistics of Challenge-MT
We compare the complete WMT test set and the Challenge-MT subset in terms of the length of source sentences, the length of target sentences, the perplexity of source sentences, average number of nouns, verbs and named entities in the source sentence. The statistics is shown in Table 4.
Appendix B Details of Experiments
We conduct experiments under the few-shot setting. To obtain the demonstrations of CoT, we ask the LLM to output the step-by-step translation process in a manner of post-explanation (i.e.,, given the source sentence and its translation, requesting the LLM to generate the intermediate process). To obtain the ones of MAPS, we let the LLM to perform translation with the specific strategy on the validation set, and assemble the generated intermediate process (e.g., keywords) and the reference translation as demonstrations.
Appendix C Results under the Rerank setting
We follow He et al. (2023) to conduct experiment additionally under the rerank setting. For the baseline ICL, we run for 4 times with different sets of demonstrations, which are sampled randomly with seeds , and adopt QE to select the best candidate as the final translation. For MAPS, the final translation is selected from the candidates generated by the three strategies (‘+topic’, ‘+Keywords’, and ‘+SimDems’) and ICL (seed=1). For xIoD, we select the final translation from the results of xIoD and ICL (seed=1). The results are shown in Table 5.
| Methods | EnZh | ZhEn | EnEt | EtEn | EnIs | IsEn | Average | |||||||
| COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | |
| Baselines | ||||||||||||||
| ICL | 76.79 | 3.94 | 72.67 | 0.43 | 82.10 | 9.37 | 79.98 | 7.44 | 73.42 | -1.21 | 78.88 | 4.80 | 77.31 | 4.13 |
| MAPS | 77.24 | 4.56 | 73.17 | 1.70 | 83.05 | 10.57 | 80.12 | 8.28 | 75.67 | 2.61 | 78.47 | 5.13 | 77.95 | 5.48 |
| Ours | ||||||||||||||
| xIoD-I | 77.36 | 4.37 | 73.30 | 1.08 | 83.06 | 10.39 | 80.22 | 7.96 | 76.88 | 3.12 | 78.93 | 5.29 | 78.29 | 5.37 |
| xIoD-E | 77.78 | 5.04 | 73.36 | 0.88 | 83.21 | 10.93 | 80.10 | 8.06 | 77.39 | 3.97 | 79.22 | 5.31 | 78.51 | 5.70 |