### v2

# Aligning Translation-Specific Understanding to General Understanding in Large Language Models

Abstract
Large Language models (LLMs) have exhibited remarkable abilities in understanding complex texts, offering a promising path towards human-like translation performance. However, this study reveals the misalignment between the translation-specific understanding and the general understanding inside LLMs. This understanding misalignment leads to LLMs mistakenly or literally translating some complicated concepts that they accurately comprehend in the general scenarios (e.g., QA). To align the translation-specific understanding to the general one, we propose a novel translation process, DUAT (Difficult words Understanding Aligned Translation), explicitly incorporating the general understanding on the complicated content incurring inconsistent understanding to guide the translation. Specifically, DUAT performs cross-lingual interpretation for the difficult-to-translate words and enhances the translation with the generated interpretations. Furthermore, we reframe the external tools to improve DUAT in detecting difficult words and generating helpful interpretations. We conduct experiments on the self-constructed benchmark Challenge-WMT111The dataset is available at: ChallengeWMT 🖂 means corresponding author., consisting of samples that are prone to mistranslation. Human evaluation results on high-resource and low-resource language pairs indicate that DUAT significantly facilitates the understanding alignment, which improves the translation quality (up to +3.85 COMET) and reduces the literality of the translation by - 25% - 51%.
UTF8gbsn
Aligning Translation-Specific Understanding to General Understanding in Large Language Models
Yichong Huang†, Baohang Li†, Xiaocheng Feng†‡, Wenshuai Huo†‡, Chengpeng Fu†‡, Ting Liu†, Bing Qin †Harbin Institute of Technology ‡ Peng Cheng Laboratory {ychuang,xcfeng,baohangli,cpfu,wshuo,tliu,qinb}@ir.hit.edu.cn
1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable language understanding and generation, paving the way for a higher level of performance in machine translation Zhao et al. (2023); OpenAI (2023); Jiang et al. (2023); Workshop (2023). However, existing research reports that LLMs have yet to achieve as significant advances in machine translation as they have achieved in other natural language processing fields Hendy et al. (2023); Pang et al. (2024); Zhang et al. (2023a); Jiao et al. (2023); Zhu et al. (2023b); Lou et al. (2023).
In this study, we discover the misalignment between the general understanding and translation-specific understanding inside LLMs, as illustrated in Fig.1. This understanding misalignment leads to LLMs mistakenly or literally translating some complicated concepts that they accurately comprehend in general scenarios. We refer to these failures as language models’ generalization failures on translation. Human evaluation on a total of 600 sampled sentences across six language pairs show that generalization failures account for a considerable proportion of all mistranslations (16%-32%), indicating serious understanding misalignment (§6.1).
To align the translation-specific understanding to the general one, we propose a novel translation process, DUAT (Difficult words Understanding Aligned Translation), explicitly incorporating the general understanding on the complicated content incurring inconsistent understanding to guide the translation. Specifically, DUAT first detects the difficult-to-translate words in the source sentence, which could cover the generalization failures intuitively. Next, the LLM is prompted to interpret each difficult word with the target language, i.e., cross-lingual interpretation, unleashing the powerful general understanding and transforming this understanding into the target language space. After that, DUAT conducts translation under the guidance of these interpretations. Unlike the CoT-based process mimicking junior translators to sequentially translate all words Peng et al. (2023), DUAT works like senior translators to analyze the complicate words, which helps the model deep understand the source sentence and produces more nuanced translations. Furthermore, we reframe the external tool of token-level QE Rei et al. (2023) to enhance the detection of difficult words, and design a strategy of interpretation quality control to filter hallucinated interpretations based on sentence-level QE Rei et al. (2020).
To better analyze the understanding misalignment, we proposed the Challenge-WMT benchmark, which contains more sentences prone to mistranslation. These sentences were collected from multi-year WMT datasets and represent difficult samples that multiple state-of-the-art (SOTA) systems translate poorly. Human evaluation results indicate that DUAT significantly facilitates the understanding alignment, reducing 80%88% of generalization failures. Moreover, this alignment improves the translation quality, as evidenced by automatic metrics (up to +3.85 COMET), and alleviates translation literalness by -25% -51%.
2 Background
2.1 LLM-based MT
Considering the translation from source language to target language , LLM-based machine translation converts the source sentence to an instruction using a translation-specific template and generates the translation by feeding the instruction to the LLM . To make the LLM better follow the instruction, the in-context learning (ICL) strategy Brown et al. (2020); Dong et al. (2023) injects a few examples/demonstrations of translation into the instruction, which is shown as:
Formally, the LLM-based MT generates the translation with ICL as:
where is the demonstrations set of translation.
2.2 Quality estimation (QE)
QE for machine translation, i.e., reference-free MT evaluation, aims to predict the quality of the given translation only according to the source sentence, which has shown auspicious correlations with human judgments Rei et al. (2020, 2021). Given a source sentence and a translation , QE score is denoted as .
Thanks to the recent advance in the interpretability of neural MT metrics Rei et al. (2023), token-level QE is proposed to score the error degree of the given translation span by calculating the misalignment of this span against the source sentence. Given a source sentence and the candidate translation , token-level QE annotates the error degree of the specific span in the translation, i.e., where .
3 Approach: DUAT
In this section, we first introduce our translation framework DUAT (§3.1). Specifically, DUAT consists of three components: difficult word detection (§3.2), cross-lingual interpretation (§3.3), and interpretation quality control (§3.4). To make the LLM follow the procedure of each component as expected, we adopt the in-context learning strategy and design an automatic method for constructing demonstrations of DUAT (§3.5).
3.1 Framework
The progress of DUAT is illustrated in Fig.2. Given the source sentence, DUAT first detects the difficult words or phrases in the source sentence. Once the difficult words are identified, DUAT requests the LLM to interpret each difficult word with the target language, unleashing the powerful understanding capability inside the LLM and transforming these understandings into the target language space. Finally, to avoid the interference of incorrect and useless interpretations, DUAT removes the negative interpretations through the interpretation quality control and outputs the final translation guided by the helpful interpretations.
3.2 Step-1: Difficult Word Detection
In practice, we found that directly inquiring LLMs to identify difficult-to-translate words is challenging. To tackle this challenge, we first conduct a preliminary translation for the given source sentence and then extract the mistranslated words and phrases in the source sentence as the difficult words. Concretely, we invent DUAT-I to do this leveraging the Intrinsic ability of LLMs at first.
DUAT-I.
Given source sentence , DUAT-I first obtains the preliminary translation (also known as draft translation) by prompting the LLM to translate with the in-context learning strategy, which is shown in Eq. (1). Next, the LLM is requested to output the difficult words based on the source sentence and the preliminary translation:
DUAT-I obtains the difficult word list via performing greedy decoding on the LLM:
where is the LLM, which is prompted with demonstrations of difficult word detection .
DUAT-E.
It is frequently observed that the LLM fails to recognize the mistranslated words, due to their limitation in self-knowledge Yin et al. (2023) Therefore, we devise DUAT-E to boost the detection with the External tool. First, DUAT-E requests the LLM with the same prompt as DUAT-I while performing temperature sampling for times. Next, the union of all sampling results is taken as the candidate set of difficult words :
where is the hyperparameter of sampling temperature, which is set to to capture more candidates, and is set to 5 empirically.
Finally, DUAT-E annotates each candidate word with its degree of misalignment with respect to the draft translation, which reflects the translation-specific difficulty. To implement this function, we adopt an external tool of token-level QE . As shown in §2.2, token-level QE is originally used to score the mistranslation degree of the given translation span with respect to the source sentence, i.e., where . Differently, we utilize this tool in a dual manner. That is, we use to annotate the misalignment degree of the given source span with respect to the translation, i.e., where . Formally, the misalignment score of each difficult word candidate is calculated as:
Then, DUAT-E selects candidates with misalignment score , where is the hyperparameter named the difficulty threshold. We refer to this procedure as the difficulty-aware selection in Fig.2.
3.3 Step-2: Cross-Lingual Interpretation
After the difficult words in the source sentence are detected, DUAT lets the LLM generate the interpretation of each difficult word via requesting:
Through access to the LLM, the interpretation set is obtained:
where , which is the demonstrations of the cross-lingual interpretation.
Prob and cons.
Under the guidance of generated interpretations, DUAT can align the translation-specific understanding to the general one. However, LLMs may generate hallucinated interpretations sometimes (e.g., the interpretation of "崩塌" in Fig.2), which biases the resulting translation from the original semantics. Besides, helpless interpretations that can not provide useful information also pose a risk of disturbing the translation process.
3.4 Step-3: Interpretation Quality Control
To overcome the potential interference of the generated negative interpretations, DUAT removes them through the interpretation quality control (IQC) and outputs the final translation guided by the helpful interpretations.
Concretely, given a set of interpretations , DUAT ablates each interpretation sequentially and uses the remaining interpretations to guide the translation. The interpretation-guided translation is implemented in a fashion of refinement:
Formally, the translation is obtained as:
where , which is the demonstration set of interpretation-guide translation.
If the better translation performance is achieved by ablation, which is measured by the QE222We use wmt21-comet-qe-da as the QE scorer. tool due to the unavailable access to the reference translation, the interpretation is removed from and the current translation is taken as the best translation. We also detail this process in Alg.1.
3.5 Demonstrations Synthesis for DUAT
To make the LLM follow the procedure of DUAT as expected, we adopt the ICL strategy. Common practice constructs demonstrations manually, necessitating human translators proficient in language pairs for languages. To overcome this considerable cost, we devise a method for synthesizing high-quality demonstrations of DUAT based on parallel data. This process of synthesizing demonstrations is accomplished in a manner of post-explanation by asking the LLM to compare the baseline translation and the reference translation. We describe this process in Appendix B.2.
Methods | EnZh | ZhEn | EnEt | EtEn | EnIs | IsEn | Average | |||||||
COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | |
Existing Systems | ||||||||||||||
74.85 | 54.95 | 68.21 | 52.65 | 79.11 | 68.71 | 78.83 | 65.46 | 76.17 | 59.67 | 78.70 | 66.54 | 75.98 | 61.33 | |
NLLB | 68.77 | 47.77 | 60.09 | 45.14 | 74.20 | 63.13 | 74.35 | 60.87 | 69.37 | 52.08 | 72.55 | 59.66 | 69.89 | 54.78 |
GPT-4 | 76.15 | 55.37 | 70.77 | 58.85 | 80.25 | 70.80 | 77.83 | 65.48 | 77.33 | 61.15 | 79.39 | 68.40 | 76.95 | 63.34 |
Baselines | ||||||||||||||
Zero-shot | 74.89 | 54.25 | 71.27 | 58.24 | 80.67 | 69.10 | 74.93 | 61.75 | 71.17 | 53.12 | 76.22 | 64.05 | 74.86 | 60.09 |
ICL | 75.47 | 55.79 | 72.22 | 59.56 | 80.9 | 69.93 | 79.40 | 66.63 | 73.19 | 54.87 | 77.52 | 65.81 | 76.45 | 62.10 |
+CoT | 73.85 | 53.42 | 71.35 | 57.90 | 78.03 | 66.03 | 76.78 | 63.97 | 69.72 | 50.98 | 76.55 | 64.54 | 74.38 | 59.47 |
+Keywords | 73.93 | 55.10 | 71.22 | 59.18 | 78.63 | 70.01 | 77.79 | 66.30 | 70.33 | 54.31 | 74.55 | 64.40 | 74.41 | 61.55 |
+Topic | 75.83 | 53.60 | 72.46 | 57.98 | 80.98 | 66.78 | 79.20 | 64.64 | 72.77 | 51.98 | 76.49 | 62.06 | 76.29 | 59.51 |
+SimDems | 75.22 | 55.10 | 72.20 | 59.16 | 81.24 | 70.42 | 79.11 | 66.29 | 72.70 | 54.04 | 76.78 | 64.43 | 76.21 | 61.57 |
Ours | ||||||||||||||
DUAT-I | 76.92 | 56.16 | 72.94 | 59.94 | 82.92 | 72.19 | 79.96 | 67.05 | 76.64 | 57.71 | 78.45 | 66.80 | 77.97 | 63.31 |
DUAT-E | 77.57 | 56.86 | 73.25 | 60.16 | 83.07 | 72.30 | 80.01 | 66.91 | 77.04 | 57.38 | 78.70 | 66.93 | 78.27 | 63.42 |
4 Testbed: Challenge-WMT
To better analyze the understanding misalignment problem of LLMs, we propose the benchmark Challenge-WMT, which contains challenge sentences that are prone to mistranslation. This benchmark is constructed by collecting samples that multiple SOTA systems translate poorly from multi-year WMT testsets of six language pairs (Chinese (zh), Estonian (et), and Icelandic (is) to/from English (en)). Additionally, we believe that this dataset could promote future research in understanding the limitations of existing MT systems.
We select three SOTA MT systems: Google Translate, ChatGPT, and NLLB NLLB Team et al. (2024). Due to the poor performance of NLLB in the zhen translation, we additionally train a zhen translation model based on DeltaLM Ma et al. (2021) on the parallel corpus from OPUS333https://opus.nlpl.eu/. Next, all of the system translations are scored with COMET metric, and the of samples with the lowest score for each system are extracted as its difficult samples set. We vary the value of across different language pairs to ensure an appropriate scale for each difficult sample set. Finally, the intersection of all systems’ difficult sample sets is taken as the Challenge-WMT testbed. Challenge-WMTcomprises around 600+ sentence pairs for each language pair, which is illustrated in Tab.6. We equally split this dataset into the validation set and the test set.
We report the translation performance measured by COMET on Challenge-WMT and the complete WMT set in Fig.8, which shows that the performance decreases dramatically on Challenge-WMT (84.573.6 averagely). Next, we conduct a multi-aspect comparison for Challenge-WMT and the complete set in Appendix C, and find that the samples of Challenge-WMT have higher perplexity (214252 averagely). This result indicates that sentences in Challenge-WMT are more complex.
5 Experiments
5.1 Experimental Setup
Comparative Methods.
We verify the effectiveness of our DUAT on the LLM GPT-3.5-turbo for its promising capability in following complicated instructions. Demonstrations of DUAT are gained by performing our automatic method (§3.5) on the validation set of Challenge-WMT. We compare DUAT with the following methods:
Zero-shot, which asks the LLM to translate the source sentence directly.
ICL (In-Context Learning), enhancing the translation with randomly selected exemplars from the validation set.
CoT Wei et al. (2022), encouraging the LLM to resolve the problem step by step. In this work, we re-implement CoT by prompting the LLM to translate the source sentence step by step.
MAPS He et al. (2024), incorporating the knowledge of keywords, topic words, and demonstrations similar to the given source sentence to enhance the translation process, respectively.
Commercial and open-source systems. We also report the performance of Google Translate, NLLB (in zhen translation, we replace NLLB with our trained MT model based on DeltaLM), and zero-shot translation based on GPT4 (GPT-4-turbo).
For DUAT and other ICL-based methods, we select =8 demonstrations (i.e., 8-shot) to achieve a strong baseline performance. More details of re-implementing the baselines under the few-shot setting are illustrated in Appendix D.
Metrics.
5.2 Results on Challenge-WMT
Methods | EnIs | IsEn | ||
COMET | BLEURT | COMET | BLEURT | |
Zero-shot | 77.33 | 61.15 | 79.39 | 68.40 |
ICL | 80.1 | 61.99 | 81.02 | 70.20 |
DUAT-E | 81.7 | 64.01 | 81.21 | 69.22 |
The main results are illustrated in Tab.1. From the results, we have drawn the following observations:
(1) DUAT achieves significant improvements.
On average, DUAT-E surpasses the baseline ICL by +1.82 COMET and +1.32 BLEURT, and improves Zero-shot by +3.41 COMET and +3.33 BLEURT. In the low-resource translation of EnIs, DUAT-E improves ICL by +3.85 COMET and Zero-shot by +5.87 COMET. These improvements show that, DUAT largely elicits the translation ability via aligning the translation-specific understanding to the general one.
(2) DUAT achieves state-of-the-art performance on Challenge-WMT.
DUAT achieves the highest scores in EnZh and EnEt translation in terms of COMET and BLEURT. In EnIs translation, GPT-4 achieves the best results due to its superior multilingual capabilities. To verify the effectiveness of DUAT on LLMs with stronger multilingual capabilities, we implement DUAT based on GPT-4 in EnIs translation, as shown in Table 2. The results suggest that DUAT can also benefit stronger multilingual LLMs by facilitating the understanding misalignment.
Methods | EnZh | ZhEn | ||
COMET | BLEURT | COMET | BLEURT | |
WMT22 Best | 86.80 | – | 81.00 | – |
[…]
ICL | 87.10 | 73.15 | 82.71 | 71.38 |
DUAT-E | 87.60 | 72.41 | 82.75 | 71.75 |
(3) CoT works poorly in machine translation.
In Table 1, CoT incurs a dramatic performance drop over the baseline ICL. Our case studies reveal that CoT produces extremely wordy translations, which is also observed by Peng et al. (2023). We conjecture that CoT makes LLMs imitate junior translators rather than advanced translators.
(4) Difficult words are the bottleneck in translating complex sentences.
The results show that incorporating the analysis of keywords and topics He et al. (2024) has yet to gain significant improvements as DUAT. It suggests that it is the difficult words that lead to the performance bottleneck in translating intricate sentences. We also follow He et al. (2024) to experiment under the rerank setting as shown in Appendix.E, which shows the effectiveness of our method further.
(5) The external tool helps LLMs better identify difficult words.
DUAT-I achieves an average improvement of +1.52 COMET, demonstrating the effective performance of LLMs in recognizing difficult words. And DUAT-E gains a further improvement of +0.3 COMET, showing the effectiveness of the external token-level QE tool in this task.
5.3 Results on the complete WMT
Experimental results on Challenge-WMT show the effectiveness of DUAT in translating complex sentences, raising the question of its impact on translating simple sentences. Therefore, we conduct additional experiments on the complete WMT2022 testset of ZhEn translation. As the results shown in Tab.5.2, our method achieves comparable results to the baseline ICL in both translation directions. These results show that DUAT has no negative impact on translating simple sentences.
EnZh | ZhEn | EnEt | EtEn | EnIs | IsEn | |
#Mistranslation | 19 | 25 | 26 | 53 | 22 | 34 |
#Generalization Failure of Baseline | 5 (26%) | 4 (16%) | 7 (27%) | 17 (32%) | 5 (23%) | 10 (29%) |
[…]
6 Analysis
6.1 Human Evaluation
To quantitatively analyze the understanding misalignment problem, we employ one senior human translator for each language direction to assess generalization failures and translation literalness. Specifically, in each direction, we randomly sample 100 sentences444It takes 1.4 dollars for annotating one sentence. from Challenge-WMT and ask the senior translator to annotate the mistranslated words and phrases (i.e., mistranslation) and score the literalness (1 to 5 score) of the translations generated by the strong baseline (GPT-3.5-turbo with ICL). Next, we ask the LLM to generate interpretations of the mistranslated content. These interpretations are presented to the translator, who judges whether they contain the correct understanding of the content. If the interpretation is accurate, the content is annotated as a generalization failure. Finally, the translator is asked to annotate the generalization failures of DUATand score the literalness of the translations produced by DUAT.
Analysis of generalization failures.
As the results shown in Tab.5.3, generalization failures account for a considerable proportion of all mistranslations (16%32%). Our method (DUAT-E) largely resolves these failures by 71%88%. We further study the unresolved failures and find that most of these unresolved failures are because they are not identified as difficult words by the LLMs in the stage of difficult word detection (Eq.3) despite actually hard to translate. It indicates that the difficult word detection remains an open question.
Analysis of translation literalness.
As the results shown in Tab.5.3, the baseline translations are highly biased towards literal translation. DUATsignificantly reduces the bias towards literal translation, indicating that the process of interpreting the difficult words first and then translating aligns better with sense-for-sense translation.
6.2 Ablation Study
Methods | EnZh | ZhEn | ||
COMET | COMET | |||
DUAT-E | 77.57 | – | 73.23 | – |
w/o. Draft | 76.94 | -0.63 | 72.68 | -0.55 |
w/o. IQC | 76.54 | -1.03 | 72.91 | -0.32 |
DUAT-I | 76.92 | – | 72.94 | – |
w/o. Draft | 76.68 | -0.24 | 72.78 | -0.16 |
w/o. IQC | 76.45 | -0.47 | 72.59 | -0.35 |
DUAT introduces the processes of (1) draft translation to precisely detect the difficult words and (2) IQC to improves the helpfulness of interpretations. To clearly elucidate the contribution of these two components, we conduct an ablation study in Table 5. Specifically, we analyze the effect of the draft translation by asking the LLM to detect difficult words directly without the draft translation. The impact of IQC is analyzed by evaluating the performance of the generated translations guided by the original noisy interpretations (i.e., without the processing of IQC). The results show that removing either component leads to performance drops, and IQC plays a more important role in DUAT. Specifically, the improvement of DUAT is halved when ablating the IQC on the EnZh translation.
6.3 Analysis of Difficult Word Detection
To offer an in-depth insight into the process of difficult word detection, we illustrate the relation between the number of difficult words interpreted and the resulting performance by adjusting the value of the difficulty threshold (), which is shown in Fig.3. Concretely, a smaller value of allows more difficult words to be interpreted. From the results, we have the following observations:
Increasing the number of interpretations does not necessarily lead to performance improvements, but increasing high-quality ones can.
Specifically, without controlling the quality of the interpretations (i.e., w/o. IQC), increasing the number of interpretations (the green lines) yields unpredictable performance changes (as shown by the green bins), as introducing either valuable information or noise. Fortunately, with IQC filtering negative interpretations, increasing the number of interpretations (the blue lines) leads to constant improvements (as the blue bins show).
Interpreting words that are more difficult brings larger improvements.
Specifically, in the EnZh translation, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to (). However, decreasing the value of from to , the average number of helpful interpretations is increased from to (), and the performance is increased from to (). It should be noted that interpreting more words incurs more inference costs. Therefore, a modest value of (i.e., ) is recommended to reach a compromise between efficiency and performance of DUAT.
6.4 Analysis of Interpretation Generation
Languages of interpretations.
Given a difficult word, DUAT generates the corresponding interpretation with the target language (i.e., cross-lingual interpretation), which implicitly comprises two stages: (1) generating the interpretation in the source language and (2) translating the interpretation into the target language. Compared with conducting these two stages explicitly, DUAT is more efficient and avoids error accumulation, which is illustrated in Fig.4. As demonstrated, interpretations in the target language (the blue bins) are more beneficial than the ones in the source language (the purple bins) owing to aligning the general understanding into the target language space, which could provide more benefits for translation. And the implicit two-stage process (the blue bins) is better than the explicit one (the green bins).
7 Related Work
Evaluation of LLMs’ translation capabilities.
With the remarkable progress of LLMs, researchers have assessed their translation abilities in various aspects. Zhang et al. (2023a); Vilar et al. (2023); Garcia et al. (2023); Bawden and Yvon (2023) first investigate LLM-based MT in terms of the prompt template and examples selection. Next, the evaluation is extended across more domains Hendy et al. (2023), more languages Zhu et al. (2023a), and document-level translation Hendy et al. (2023); Wang et al. (2023). Other lines of work have performed in-depth assessments on the important attributes beyond accuracy, like literalness Raunak et al. (2023) and culture awareness Yao et al. (2023). As existing studies have shown that LLMs have achieved promising performance, our work turns out to benchmark them on hard instances towards detecting more underlying issues.
LLM-based translation strategies.
Lu et al. (2023) obtain the multilingual translations of keywords in the source sentence via the translator NLLB to augment the LLM, which improves the translation of low-resource languages while hurting the performance of high-source languages. Chen et al. (2023) demonstrate that iterative refinement reduces translationese significantly. He et al. (2024) incorporate the knowledge of keywords, topics, and reference demonstrations to enhance the translation process, and use a rerank strategy to combine all candidate translations. However, there is no significant improvement to be observed when solely utilizing each single type of knowledge. Different from previous works that utilize the intrinsic knowledge of LLM, DUAT focuses on dealing with the difficult-to-translate words instead of the keywords for the reason that we argue the difficult-to-translate words lead to the performance bottleneck due to the long-tail distribution of knowledge.
LLM-based Automatic Post-Editing (APE).
APE corrects the errors in the generated translation, aiming to bias the translation towards the distribution of the target language (Chen et al., 2023; Koneru et al., 2024). Differently, our work aims to leverage the powerful understanding abilities of LLMs to correct the misunderstanding of complicated concepts in the source sentence. Our work is in parallel with APE.
8 Conclusion
In this work, we propose a novel translation process, DUAT, to take the first step in resolving the misalignment between the translation-specific understanding and the general understanding. Furthermore, we utilize the token-level QE tool to enhance the detection of difficult words and the sentence-level QE tool to remove harmful interpretations. Human evaluation results on high-resource and low-resource language pairs indicate that DUAT significantly facilitates the understanding alignment, which improves the translation quality and alleviates translation literalness.
9 Limitations
Even though DUAT elicits the translation abilities of LLMs via unleashing the general understanding (intrinsic knowledge) of LLMs, they still struggle to translate concepts that require the incorporation of extrinsic knowledge, such as the translation of neologisms. However, Our approach lays the foundation for researching when and how to incorporate external knowledge. Besides, DUAT requires to prompt the LLM for several times, leading to an increase in latency. This latency is mainly caused by our interpretation quality control (IQC) strategy, which sequentially ablates each generated interpretation. Concretely, if difficult words are identified, IQC needs to prompt the LLM for times.
Acknowledgements
Bing Qin is the corresponding author of this work. We thank the anonymous reviewers for their insightful comments. This work was supported by the National Natural Science Foundation of China (NSFC) (U22B2059, grant62276078), the Key R&D Program of Heilongjiang via grant 2022ZX01A32, the International Cooperation Project of PCL, PCL2022D01, the Nature Scientific Foundation of Heilongjiang Province(YQ2021F006), and the Fundamental Research Funds for the Central Universities (Grant No.HIT.OCEF.2023018).
Appendix A Generalization Failures on Translation
In this section, we first provide the illustration of generalization failures on more LLMs, as shown in Fig.5. As we can see, all of four LLMs accurately comprehend the complex concept, which three out of them mistranslate this concept. Then, we also provide more examples of generalization failures, as shown in Fig.7 and Fig.7.
Appendix B More details of DUAT
B.1 Details of IQC
We gives a formal description of our interpretation quality control in Alg.1.
B.2 Details of Demonstration Synthesis
Inspired by the idea of Auto-CoT Zhang et al. (2023b), we utilize LLM to generate the difficult words and corresponding interpretations based on the given bilingual sentence pair :
Then, the response is parsed via regular expression to extract the difficult words and interpretations . Next, we remove the noisy interpretations through a process similar to IQC (Alg. 1). The only difference is that the QE metric is replaced with the reference-based COMET Rei et al. (2020) due to the available access to the reference translation. Finally, the generated difficult words and interpretations can be assembled with the source and target sentence as demonstrations for each step of DUAT.
Appendix C Statistics of Challenge-WMT
We compare the complete WMT test set and the Challenge-WMT subset in terms of the length of source sentences, the length of target sentences, the perplexity of source sentences, average number of nouns, verbs and named entities in the source sentence. The statistics is shown in Table 6.
Appendix D Details of Experiments
We conduct experiments under the few-shot setting. To obtain the demonstrations of CoT, we ask the LLM to output the step-by-step translation process in a manner of post-explanation (i.e.,, given the source sentence and its translation, requesting the LLM to generate the intermediate process). To obtain the ones of MAPS, we let the LLM to perform translation with the specific strategy on the validation set, and assemble the generated intermediate process (e.g., keywords) and the reference translation as demonstrations.
Appendix E Results under the Rerank setting
We follow He et al. (2024) to conduct experiments additionally under the rerank setting. For the baseline ICL, we run for 4 times with different sets of demonstrations, which are sampled randomly with seeds , and adopt QE to select the best candidate as the final translation. For MAPS, the final translation is selected from the candidates generated by the three strategies (‘+topic’, ‘+Keywords’, and ‘+SimDems’) and ICL (seed=1). For DUAT, we select the final translation from the results of DUAT and ICL (seed=1). The results are shown in Table 7.
Methods | EnZh | ZhEn | EnEt | EtEn | EnIs | IsEn | Average | |||||||
COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | COMET | QE | |
Baselines | ||||||||||||||
ICL | 76.79 | 3.94 | 72.67 | 0.43 | 82.10 | 9.37 | 79.98 | 7.44 | 73.42 | -1.21 | 78.88 | 4.80 | 77.31 | 4.13 |
MAPS | 77.24 | 4.56 | 73.17 | 1.70 | 83.05 | 10.57 | 80.12 | 8.28 | 75.67 | 2.61 | 78.47 | 5.13 | 77.95 | 5.48 |
Ours | ||||||||||||||
DUAT-I | 77.36 | 4.37 | 73.30 | 1.08 | 83.06 | 10.39 | 80.22 | 7.96 | 76.88 | 3.12 | 78.93 | 5.29 | 78.29 | 5.37 |
DUAT-E | 77.78 | 5.04 | 73.36 | 0.88 | 83.21 | 10.93 | 80.10 | 8.06 | 77.39 | 3.97 | 79.22 | 5.31 | 78.51 | 5.70 |